OpenStudy (anonymous):

man climbs 1576 steps in 9minutes and 33seconds. IF the height gain per step was 0.20m, and the mass of the man was 70.0kg, what was his average power output?

OpenStudy (anonymous):

You have to find out work first: W=Fxd F = 70 x 10 = 700N d = 0.2x1576=315.2m so W=220640 J 9 minutes 33 seconds = 573 seconds P = W/t P=220640/573 P=385.06 I'm not entirely sure though, I only learned this recently, haha

OpenStudy (anonymous):

Yep, that's right. I had actually figured it out, but thanks for the help anyway.